# A Path to Constraints on Common Envelope Ejection in Massive Binaries: Full Evolutionary Reconstruction of Three Black Hole X-ray Binaries

### 12 Apr 2026 v1

@@ V Conclusions
We thank the referee for a very careful reading and constructive comments that have led to the improvement of the manuscript. This work is supported by the Natural Science Foundation of China (grant Nos. 12125303, 12525304, 12288102, 12473034, 12090040/3, 12273105, 11703081, 11422324, 12073070, 12173081), the CAS Project for Young Scientists in Basic Research (YSBR-148), the Strategic Priority Research Program of the Chinese Academy of Sciences (grant Nos. XDB1160303, XDB1160201, XDB1160000), the National Key R&\&D Program of China (grant Nos. 2021YFA1600403, 2021YFA1600400), the Yunnan Revitalization Talent Support Program-Science &\& Technology Champion Project (No. 202305AB350003) and Young Talent project, the International Centre of Supernovae (ICESUN), Yunnan Key Laboratory of Supernova Research (Nos. 202302AN360001, 202201BC070003), Yunnan Fundamental Research Projects (No. 202401AT070139), the Natural Science Foundation of Henan Province (No. 242300420944). X.C. acknowledges the New Cornerstone Science Foundation through the XPLORER PRIZE. The authors gratefully acknowledge the “PHOENIX Supercomputing Platform” jointly operated by the Binary Population Synthesis Group and the Stellar Astrophysics Group at Yunnan Observatories, Chinese Academy of Sciences.

### 26 Apr 2026 v2

@@ V Conclusions
We deeply thank the referee for a very careful reading and constructive comments that have led to the improvement of the manuscript. The authors are grateful to Poshak Gandhi for his valuable suggestions and feedback on this work. This work is supported by the Natural Science Foundation of China (grant Nos. 12125303, 12525304, 12288102, 12473034, 12090040/3, 12273105, 11703081, 11422324, 12073070, 12173081), the CAS Project for Young Scientists in Basic Research (YSBR-148), the Strategic Priority Research Program of the Chinese Academy of Sciences (grant Nos. XDB1160303, XDB1160201, XDB1160000), the National Key R&\&D Program of China (grant Nos. 2021YFA1600403, 2021YFA1600400), the Yunnan Revitalization Talent Support Program-Science &\& Technology Champion Project (No. 202305AB350003) and Young Talent project, the International Centre of Supernovae (ICESUN), Yunnan Key Laboratory of Supernova Research (Nos. 202302AN360001, 202201BC070003), Yunnan Fundamental Research Projects (No. 202401AT070139), the Natural Science Foundation of Henan Province (No. 242300420944). X.C. acknowledges the New Cornerstone Science Foundation through the XPLORER PRIZE. The authors gratefully acknowledge the “PHOENIX Supercomputing Platform” jointly operated by the Binary Population Synthesis Group and the Stellar Astrophysics Group at Yunnan Observatories, Chinese Academy of Sciences.
Example MESA inlists we used to generate our models are archived on Zenodo and can be downloaded at 10.5281/zenodo.19660373 (catalog doi:10.5281/zenodo.19660373).